# Gradient-based Sample Selection for Faster Bayesian Optimization

### v2

@@ Abstract
Bayesian optimization (BO) is an effective technique for black-box optimization. However, its applicability is typically limited to moderate-budget problems due to the cubic complexity in computing the Gaussian process (GP) surrogate model. In large-budget scenarios, directly employing the standard GP model faces significant challenges in computational time and resource requirements. In this paper, we propose a novel approach, gradient-based sample selection Bayesian Optimization (GSSBO), to enhance the computational efficiency of BO. The GP model is constructed on a selected set of samples instead of the whole dataset. These samples are selected by leveraging gradient information to remove redundancy while preserving diversity and representativeness. We provide a theoretical analysis of the gradient-based sample selection strategy and obtain explicit sublinear regret bounds for our proposed framework. Extensive experiments on synthetic and real-world tasks demonstrate that our approach significantly reduces the computational cost of GP fitting in BO while maintaining optimization performance comparable to baseline methods.
@@ 1 Introduction
Bayesian optimization (BO) is a derivative-free optimization method for black-box functions [1, 2, 3] that enjoys enormous success in many applications such as machine learning [4], structural design [5], robotics [6], additive manufacturing [7], inertial confinement fusion design  [8], etc. BO’s strength lies in its ability to represent the unknown objective function through a surrogate model and by optimizing an acquisition function [2, 3]. BO consists of a surrogate model, which provides a global prediction for the unknown objective function, and an acquisition function that serves as a criterion to determine the next sample to evaluate. In particular, the Gaussian process (GP) model is often preferred as the surrogate model due to its versatility and reliable uncertainty estimation. However, the GP model often suffers from large data sets, making it more suitable for small-budget scenarios [4]. To fit a GP model, the dominant complexity in computing the inversion of the covariance matrix is $\mathcal{O}(n^{3})$, where $n$ is the number of data samples. As the sample set grows, the computational burden increases substantially. This limitation poses a significant challenge for scaling BO to real-world problems with large sample sets.
@@
Figure 1: Illustration of GP fitting with sample selection. Left: GP fitted with 10 samples. Right: GP fitted with 6 selected samples. With fewer selected samples, we can still fit a good GP to estimate the black box function, which can also guide us in finding the global optimum.
@@
Despite the various approaches to improve the computational efficiency of BO, including parallel BO [5, 6, 9, 10], kernel approximation [11, 12, 13, 14] and sparse GP [15, 16, 17], the computational overhead remains a burden in practice [18]. Kernel approximation methods typically involve simplifying or approximating the kernel matrix, which can degrade accuracy and lead to suboptimal performance if the approximation is too coarse. Sparse GP methods, while reducing computational complexity, introduce additional complexity through corrective terms to maintain approximation accuracy, and existing iterative implementations in Sparse GP [17] frequently involve adding new samples while removing previous ones, potentially causing inefficiencies or suboptimal sample selection in complex optimization landscapes.
@@
During the iterative search process of BO, some samples can become redundant and contribute little to the additional information gain. Such samples collected in earlier stages thus diminish in importance as the process evolves. For instance, excessive searching around identified minima becomes redundant once the optimal value has been determined, as these samples cease to offer meaningful insights for further optimization. To efficiently fit a GP, it is essential to focus on samples that provide the most informative contributions. As shown in Figure 1, carefully selected samples can effectively fit a GP. Despite the reduced number of samples, the GP still captures the key trends and features of the true function while maintaining reasonable uncertainty bounds. In this paper, we propose incorporating the gradient-based sample selection technique into the BO framework to enhance its scalability and effectiveness in large-budget scenarios. This technique was originally proposed for continual learning with online data stream [19]. The previous data are selectively sampled and stored in a replay buffer to prevent catastrophic forgetting and enhance model fitting. By using gradient information to gauge the value of each sample, one can more judiciously decide which samples are most essential for building a subset, and maintain the most representative subset of BO samples. This subset is then used to fit the GP model, accelerating the BO process while ensuring efficient and effective GP fitting. We summarize our main contributions as follows:
@@
  • Theoretical analysis. We provide a rigorous theoretical analysis of the regret bound for the GSSBO. Theoretical results show that the regret bound of our proposed algorithm is similar to that of the standard GP-UCB algorithm.
@@ 2 Related Works
BO with Resource Challenges. In practical applications, BO faces numerous challenges, including high evaluation costs, input-switching costs, resource constraints, and high-dimensional search spaces. Researchers have proposed a variety of methods to address these issues. For instance, parallel BO employs batch sampling to improve efficiency in large-scale or highly concurrent scenarios [5, 6, 9, 10]. Kernel approximation methods, such as random Fourier features, map kernels onto lower-dimensional feature spaces, thus accelerating kernel-based approaches [20, 11]. Multi-fidelity BO leverages coarse simulations with a limited number of high-fidelity evaluations to reduce the overall cost [21]. For high-dimensional tasks, random embeddings or active subspaces help reduce the search dimensionality [22, 23]. Meanwhile, sparse GP significantly reduces computational complexity by introducing inducing points [15, 16, 17, 24]. However, these approaches face limitations in practical scenarios and usually sacrifice performance for scaling. Calandriello et al. [25] scaling GP optimization by repeatedly evaluating each selected point until its posterior uncertainty falls below a preset threshold, thus limiting the number of datasets. However, the dataset still grows with time, and the algorithm’s dependence on its initial sample set means that low-value points selected early on remain permanently in the model, potentially inflating computational overhead.
@@
BO with Gradient Information. The availability of derivative information can significantly simplify optimization problems. Ahmed et al. [26] highlights the potential of incorporating gradient information into BO methods and advocates for its integration into optimization frameworks. Wu et al. [27] introduced the parallel knowledge gradient method for batch BO, achieving faster convergence to global optima. Then they introduce d-KG [28] as a new acquisition function, which systematically introduces gradient information into the BO and proves that using gradient information can strictly improve the information value. Rana et al. [29] incorporated GP priors to enable gradient-based local optimization. Chen et al. [30] proposed a unified particle-optimization framework using Wasserstein gradient flows for scalable Bayesian sampling. Bilal et al. [31] demonstrated that BO with gradient-boosted regression trees performed well in cloud configuration tasks. Tamiya et al. [32] developed stochastic gradient line BO (SGLBO) for noise-robust quantum circuit optimization. Penubothula et al. [33] funded local critical points by querying where the predicted gradient is zero. Zhang et al. [34] introduced BO of gradient trajectory (BOGAT) for efficient imaging optimization. Makrygiorgos et al. [35] integrate exact gradient observations into the Bayesian neural network surrogate’s training loss. Although these methods leverage gradient information to improve optimization efficiency and performance, they mainly focus on refining the GP model or acquisition functions.
@@
Subset Selection. Subset selection is a key task in fields such as regression, classification, and model selection, aiming to improve efficiency by selecting a subset of features or data. Random subset selection, a simple and widely used method, involves randomly sampling data, often for cross-validation or bootstrap [36]. Importance-based selection focuses on high-value data points, while active learning targets samples that are expected to provide the most information, improving model learning [37]. Filter methods rank features using statistical measures such as correlation or variance, selecting the top-ranked ones for modeling [38]. Narendra et al. [39] introduced a branch-and-bound algorithm for efficient feature selection. Yang et al. [40] proposed dataset pruning, an optimization-based sample selection method that identifies the smallest subset of training data to reduce training costs. Ash et al. [41] employs the k-means++ algorithm in the gradient space for diversity sampling in active learning. Oglic [42]first maps each data point into the reproducing kernel Hilbert space (RKHS), then uses a max–min coverage strategy in the RKHS to sequentially sample $K$ landmarks, which are employed to construct the Nyström low-rank approximation. This method is dependent on the quality of the selected landmarks. Hayakawa [43] provides tighter expected error bounds under a continuous measure for the same underlying idea. However, computing the Mercer decomposition in high dimensions incurs substantial computational cost and suffers from severe error degradation. Because both of these methods perform their approximations in RKHS at high computational expense, they are most suitable for offline, batch-mode resampling. Zhu et al. [44] proposed a “pilot estimate” to approximate the gradient of the objective function. The core idea is to compute the gradient information corresponding to each data point based on an initial parameter estimate and identify data points with larger gradient values as more “important” samples for subsequent optimization. Despite these advancements, directly applying subset selection methods to BO often yields suboptimal results, necessitating further exploration to integrate subsampling effectively into BO frameworks.
@@ 3.1 Bayesian Optimization and Gaussian Processes
BO aims to find the global optimum $x^{*}\in\mathcal{X}$ of an unknown reward function $f:\mathcal{X}\to\mathbb{R}$, over the $n$-dimensional input space $\mathcal{X}=[0,1]^{n}$. Throughout this paper, we consider maximization problems, i.e., we aim to find $x^{*}\in\mathcal{X}$ such that $f(x^{*})\leq f(x)$ for all $x\in\mathcal{X}$, get the optimal point $x^{*}=\arg\max_{x\in\mathcal{X}}f(x)$ as quickly as possible. GPs are one of the fundamental components in BO, providing a theoretical framework for modeling and prediction in a black-box function. In each round, a sample $x_{t}$ is selected based on the current GP’s posterior and acquisition function. The observed values $y_{t}$ and $x_{t}$ are then stored in the sample buffer, and the GP surrogate is updated according to these samples. This iterative process of sampling and updating continues until the optimization objectives are achieved or the available budget is exhausted. The key advantage of GPs lies in their nonparametric nature, allowing them to model complex functions without assuming a specific form. GPs are widely used for regression (Gaussian Process Regression [45], GPR) and classification tasks due to their flexibility and ability to provide uncertainty estimates. Formally, a GP can be defined as: $f(\mathbf{x})\sim\mathcal{GP}(\mu(x),k(x,x^{\prime})))$, where $\mu(x)$ is the mean function, often assumed to be zero, and $k(x,x^{\prime})$ is the covariance function, defining the similarity between points $x$ and $x^{\prime}$. It should be noted that the algorithmic complexity of GP updates is $\mathcal{O}(n^{3})$, where $n$ is the number of observed samples. As the sample set grows, the computational resources required for these updates can become prohibitively expensive, especially in large-scale optimization problems.
@@ 3.2 Diversity-based Subset Selection
Due to limited computing resources, properly selecting samples instead of using all samples to fit a model is more efficient in problems with a large sample set. In continual learning, this helps overcome the catastrophic forgetting of previously seen data when faced with online data streams. Suppose that we have a model fitted on observed samples $\mathcal{D}\triangleq\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}\in\mathcal{X}$ and $y_{i}$ is the corresponding observation. In the context of subset selection, our objective is to ensure that each newly added sample contributes meaningfully to the optimization process. That is, a constraint ensures that when we select new samples for the sample subset, the performance of the model after the new samples are added will not be worse than the performance of the previous subset samples. Let $\mathbf{g}_{n}$ be the gradient of the sample at time $n$. Following [19], we rephrase the constraints with respect to the gradients. Specifically, the constraint can be rewritten as $\langle\mathbf{g}_{n},\mathbf{g}_{i}\rangle\geq 0,\forall i\in\{1,\dots,n-1\}.$ This transformation simplifies the constraint by focusing on the inner product of the gradients, which are nonnegative, such that there will be no performance degradation. To solve the constraint, we can use the geometric properties of the gradients in the parameter space. Note that optimizing the solid angle subtended by the gradients is computationally expensive. According to the derivation in [19], the sample selection problem is equivalent to maximizing the variance of the gradient direction of the samples in the fixed-size buffer. By maximizing the variance of the gradient directions, we ensure that the selected samples represent diverse regions of the parameter space, and therefore the buffer contains diverse samples, each contributing unique information to the optimization process. How to determine the buffer size will be detailed in Section 4.3. The previous problem thus becomes a surrogate for selecting a subset $\mathcal{U}$ of the samples that maximizes the diversity of their gradients:
@@
Here, $M$ denotes the buffer size and $/\|\textbf{g}\|$ g ∥ is the normalized gradient vector. The larger the value of this formula, the more dispersed the gradient direction of the selected samples is (the higher the diversity is). The reformulated problem (1) transforms the sample selection process from a sequential approach (adding samples to the subset one at a time) into a batch selection approach (samples are selected all at once). This empirical surrogate objective is agnostic to how gradient information is computed, making it straightforward to integrate into subset-based methods.

### v3

@@ Abstract
Bayesian optimization (BO) is an effective technique for black-box optimization. However, its applicability is typically limited to moderate-budget problems due to the cubic complexity of fitting the Gaussian process (GP) surrogate model. In large-budget scenarios, directly employing the standard GP model faces significant challenges in computational time and resource requirements. In this paper, we propose a novel approach, gradient-based sample selection Bayesian Optimization (GSSBO), to enhance the computational efficiency of BO. The GP model is constructed on a selected set of samples instead of the whole dataset. These samples are selected by leveraging gradient information to remove redundancy while preserving diversity and representativeness. We provide a theoretical analysis of the gradient-based sample selection strategy and obtain explicit sublinear regret bounds for our proposed framework. Extensive experiments on synthetic and real-world tasks demonstrate that our approach significantly reduces the computational cost of GP fitting in BO while maintaining optimization performance comparable to baseline methods.
Keywords Bayesian optimization, Large-scale, Gaussian process, Gradient information, Subset selection, Sublinear regret bound, Faster
@@ 1 Introduction
Bayesian optimization (BO) [Frazier, 2018] is a successful approach to black-box optimization that has been applied in a wide range of applications, such as hyperparameter optimization and mineral resource exploration. BO’s strength lies in its ability to represent the unknown objective function through a surrogate model and by optimizing an acquisition function [Garnett, 2023, Wang et al., 2023]. BO consists of a surrogate model, which provides a global prediction for the unknown objective function, and an acquisition function that serves as a criterion to determine the next sample to evaluate. In particular, the Gaussian process (GP) model is often preferred as the surrogate model due to its versatility and reliable uncertainty estimation. However, the GP model often suffers from large data sets, making it more suitable for small-budget scenarios [Binois and Wycoff, 2022]. To fit a GP model, the dominant complexity in computing the inversion of the covariance matrix is $\mathcal{O}(n^{3})$, where $n$ is the number of data samples. As the sample set grows, the computational burden increases substantially. This limitation poses a significant challenge for scaling BO to real-world problems with large sample sets.
@@
Figure 1: Illustration of GP fitting with sample selection. Left: GP fitted with 10 samples. Right: GP fitted with 6 selected samples. With fewer selected samples, we can still fit a good GP to estimate the black box function, guiding us in finding the global optimum.
@@
Despite the various approaches to improve the computational efficiency of BO, including parallel BO [González et al., 2016, Daulton et al., 2021, 2020, Eriksson et al., 2019], kernel approximation [Kim et al., 2021, Jimenez and Katzfuss, 2023, Hensman et al., 2013, Williams and Seeger, 2000] and sparse GP [Lawrence et al., 2002, Leibfried et al., 2020, McIntire et al., 2016], the computational overhead remains a burden in practice [Shahriari et al., 2015]. Kernel approximation methods typically involve simplifying or approximating the kernel matrix, which can degrade accuracy and lead to suboptimal performance if the approximation is too coarse. Sparse GP methods, while reducing computational complexity, introduce additional complexity through corrective terms to maintain approximation accuracy, and existing iterative implementations in Sparse GP [McIntire et al., 2016] frequently involve adding new samples while removing previous ones, potentially causing inefficiencies or suboptimal sample selection in complex optimization landscapes.
@@
During the iterative search process of BO, some samples can become redundant and contribute little to the additional information gain. Such samples collected in earlier stages thus diminish in importance as the process evolves. For instance, excessive searching around identified minima becomes redundant once the optimal value has been determined, as these samples cease to offer meaningful insights for further optimization. To efficiently fit a GP, it is essential to focus on samples that provide the most informative contributions. As shown in Figure 1, carefully selected samples can effectively fit a GP. Despite the reduced number of samples, the GP still captures the key trends and features of the true function while maintaining reasonable uncertainty bounds. In this paper, we propose incorporating the gradient-based sample selection technique into the BO framework to enhance its scalability and effectiveness in large-budget scenarios. This technique originated in continual learning with online data stream [Aljundi et al., 2019]. The previous data are selectively sampled and stored in a replay buffer to prevent catastrophic forgetting and enhance model fitting. By using gradient information to gauge the value of each sample, one can more judiciously decide which samples are most essential for building a subset, and maintain the most representative subset of BO samples. This subset is then used to fit the GP model, accelerating the BO process while ensuring efficient and effective GP fitting. To the best of our knowledge, this work is the first to propose using gradient information for subset selection to accelerate Bayesian optimization. We summarize our main contributions as follows:
@@
  • Theoretical analysis. We provide a rigorous theoretical analysis of the regret bound for the GSSBO. Theoretical results show that the regret bound of our proposed algorithm is similar to that of the standard GP-UCB.
@@ 2 Related Works
BO with Resource Challenges. In practical applications, BO faces numerous challenges, including high evaluation costs, input-switching costs, resource constraints, and high-dimensional search spaces. Researchers have proposed a variety of methods to address these issues. For instance, parallel BO employs batch sampling to improve efficiency in large-scale or highly concurrent scenarios [González et al., 2016, Daulton et al., 2021, 2020, Eriksson et al., 2019]. Kernel approximation methods, such as random Fourier features, map kernels onto lower-dimensional feature spaces, thus accelerating kernel-based approaches [Rahimi and Recht, 2007, Kim et al., 2021]. Multi-fidelity BO leverages coarse simulations with a limited number of high-fidelity evaluations to reduce the overall cost [Kandasamy et al., 2016]. For high-dimensional tasks, random embeddings or active subspaces help reduce the search dimensionality [Wang et al., 2016, Nayebi et al., 2019]. Meanwhile, sparse GP significantly reduces computational complexity by introducing inducing points [Lawrence et al., 2002, Leibfried et al., 2020, McIntire et al., 2016, Moss et al., 2023]. However, these approaches face limitations in practical scenarios and usually sacrifice performance for scaling. Calandriello et al. [Calandriello et al., 2022] scaling GP optimization by repeatedly evaluating each selected point until its posterior uncertainty falls below a preset threshold, thus limiting the number of datasets. However, the dataset still grows with time, and the algorithm’s dependence on its initial sample set means that low-value points selected early on remain permanently in the model, potentially inflating computational overhead.
@@
BO with Gradient Information. The availability of derivative information can significantly simplify optimization problems. Ahmed et al. [Ahmed et al., 2016] highlights the potential of incorporating gradient information into BO methods and advocates for its integration into optimization frameworks. Wu et al. [Wu and Frazier, 2016] introduced the parallel knowledge gradient method for batch BO, achieving faster convergence to global optima. Then they introduce d-KG [Wu et al., 2017] as a new acquisition function, which systematically introduces gradient information into the BO and proves that using gradient information can strictly improve the information value. Rana et al. [Rana et al., 2017] incorporated GP priors to enable gradient-based local optimization. Chen et al. [Chen et al., 2018] proposed a unified particle-optimization framework using Wasserstein gradient flows for scalable Bayesian sampling. Bilal et al. [Bilal et al., 2020] demonstrated that BO with gradient-boosted regression trees performed well in cloud configuration tasks. Tamiya et al. [Tamiya and Yamasaki, 2022] developed stochastic gradient line BO (SGLBO) for noise-robust quantum circuit optimization. Penubothula et al. [Penubothula et al., 2021] funded local critical points by querying where the predicted gradient is zero. Zhang et al. [Zhang and Rodgers, 2024] introduced BO of gradient trajectory (BOGAT) for efficient imaging optimization. Makrygiorgos et al. [Makrygiorgos et al., 2025] integrate exact gradient observations into the Bayesian neural network surrogate’s training loss. Although these methods leverage gradient information to improve optimization efficiency and performance, they mainly focus on refining the GP model or acquisition functions.
@@
Subset Selection. Subset selection is a key task in fields such as regression, classification, and model selection, aiming to improve efficiency by selecting a subset of features or data. Random subset selection, a simple and widely used method, involves randomly sampling data, often for cross-validation or bootstrap [Hastie, 2009]. Importance-based selection focuses on high-value data points, while active learning targets samples that are expected to provide the most information, improving model learning [Quinlan, 1986]. Filter methods rank features using statistical measures such as correlation or variance, selecting the top-ranked ones for modeling [Guyon and Elisseeff, 2003]. Narendra et al. [Narendra and Fukunaga, 1977] introduced a branch-and-bound algorithm for efficient feature selection. Yang et al. [Yang et al., 2022] proposed dataset pruning, an optimization-based sample selection method that identifies the smallest subset of training data to reduce training costs. Ash et al. [Ash et al., 2019] employs the k-means++ algorithm in the gradient space for diversity sampling in active learning. Oglic et al. [Oglic and Gärtner, 2017] first maps each data point into the reproducing kernel Hilbert space (RKHS), then uses a max–min coverage strategy in the RKHS to sequentially sample $K$ landmarks, which are employed to construct the Nyström low-rank approximation. This method is dependent on the quality of the selected landmarks. Hayakawa et al. [Hayakawa et al., 2023] provides tighter expected error bounds under a continuous measure for the same underlying idea. However, computing the Mercer decomposition in high dimensions incurs substantial computational cost and suffers from severe error degradation. Because both of these methods perform their approximations in RKHS at high computational expense, they are most suitable for offline, batch-mode resampling. Zhu et al. [Zhu, 2016] proposed a “pilot estimate” to approximate the gradient of the objective function. The core idea is to compute the gradient information corresponding to each data point based on an initial parameter estimate and identify data points with larger gradient values as more “important” samples for subsequent optimization. Despite these advancements, directly applying subset selection methods to BO often yields suboptimal results, necessitating further exploration to integrate subsampling effectively into BO frameworks.
@@ 3.1 Bayesian Optimization and Gaussian Processes
BO aims to find the global optimum $x^{*}\in\mathcal{X}$ of an unknown reward function $f:\mathcal{X}\to\mathbb{R}$, over the $n$-dimensional input space $\mathcal{X}=[0,1]^{n}$. Throughout this paper, we consider maximization problems, i.e., we aim to find $x^{*}\in\mathcal{X}$ such that $f(x^{*})\geq f(x)$ for all $x\in\mathcal{X}$, get the optimal point $x^{*}=\arg\max_{x\in\mathcal{X}}f(x)$ as quickly as possible. GPs are one of the fundamental components in BO, providing a theoretical framework for modeling and prediction in a black-box function. In each round, a sample $x_{t}$ is selected based on the current GP’s posterior and acquisition function. The observed values $y_{t}$ and $x_{t}$ are then stored in the sample buffer, and the GP surrogate is updated according to these samples. This iterative process of sampling and updating continues until the optimization objectives are achieved or the available budget is exhausted. The key advantage of GPs lies in their nonparametric nature, allowing them to model complex functions without assuming a specific form. GPs are widely used for regression (Gaussian Process Regression [Schulz et al., 2018], GPR) and classification tasks due to their flexibility and ability to provide uncertainty estimates. Formally, a GP can be defined as: $f(\mathbf{x})\sim\mathcal{GP}(\mu(x),k(x,x^{\prime}))$, where $\mu(x)$ is the mean function, often assumed to be zero, and $k(x,x^{\prime})$ is the covariance function, defining the similarity between points $x$ and $x^{\prime}$. It should be noted that the algorithmic complexity of GP updates is $\mathcal{O}(n^{3})$, where $n$ is the number of observed samples. As the sample set grows, the computational resources required for these updates can become prohibitively expensive, especially in large-scale optimization problems.
@@ 3.2 Diversity-based Subset Selection
Due to limited computing resources, properly selecting samples instead of using all samples to fit a model is more efficient in problems with a large sample set. In continual learning, this helps overcome the catastrophic forgetting of previously seen data when faced with online data streams. Suppose that we have a model fitted on observed samples $\mathcal{D}\triangleq\{(x_{1},y_{1}),\ldots,(x_{t},y_{t})\}$, where $x_{i}\in\mathcal{X}$ and $y_{i}$ is the corresponding observation. In the context of subset selection, our objective is to ensure that each newly added sample contributes meaningfully to the optimization process. That is, a constraint ensures that when we select new samples for the sample subset, the performance of the model after the new samples are added will not be worse than the performance of the previous subset samples. Let $\mathbf{g}_{t}$ be the gradient of the sample at time $t$. Following [Aljundi et al., 2019], we rephrase the constraints with respect to the gradients. Specifically, the constraint can be rewritten as $\langle\mathbf{g}_{t},\mathbf{g}_{i}\rangle\geq 0,\forall i\in\{1,\dots,n-1\}.$ This transformation simplifies the constraint by focusing on the inner product of the gradients, which are nonnegative, such that there will be no performance degradation. To solve the constraint, we can use the geometric properties of the gradients. Note that optimizing the solid angle subtended by the gradients is computationally expensive. According to the derivation in [Aljundi et al., 2019], the sample selection problem is equivalent to maximizing the variance of the gradient direction of the samples in the fixed-size buffer. By maximizing the variance of the gradient directions, we ensure that the selected samples represent diverse regions, and therefore the buffer contains diverse samples, each contributing unique information to the optimization process. How to determine the buffer size will be detailed in Section 4.3. The previous problem thus becomes a surrogate for selecting a subset $\mathcal{U}$ of the samples that maximizes the diversity of their gradients:
@@
Here, $M$ denotes the buffer size and $\textbf{g}/\|\textbf{g}\|$ is the normalized gradient vector. The larger the value of this formula, the more dispersed the gradient direction of the selected samples is (the higher the diversity is). The reformulated problem (1) transforms the sample selection process from a sequential approach (adding samples to the subset one at a time) into a batch selection approach (samples are selected all at once). This empirical surrogate objective is agnostic to how gradient information is computed, making it straightforward to integrate into subset‐based methods.